# An Empirical Study of Sensitive Information in Logs

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software safetyccs: Security and privacy Privacy protections
@@ 1. Introduction
Software logs are record statements used by software developers to capture valuable runtime information about software systems. They are generated by logging statements inserted into the source code, which produce execution logs during runtime (Barik et al., 2016; Li et al., 2020). These logs are then either appended to log files for later analysis or output to the standard console for immediate monitoring. Logging is crucial for tasks such as failure diagnosis and anomaly detection (Cinque et al., 2012; Batoun et al., 2024; Yuan et al., 2012; Aghili et al., 2023). For instance, logs are often the only resource available for diagnosing field failures (Yuan et al., 2010). By recording different events and activities within a software system, logs help developers and other software practitioners manage and maintain the health and performance of their systems effectively.
As the volume of data generated by software systems increases, the likelihood of capturing sensitive information within logs also rises, making privacy preservation more critical (Jain et al., 2016). Protecting privacy is essential not only to comply with regulatory requirements but also to maintain user trust and prevent potential harm that could arise from data leaks. A significant concern in software log management and sharing is the exposure of Personally Identifiable Information (PII). PII refers to any data that can be used to identify an individual, either directly, such as names and email addresses, or indirectly, such as device identifiers. While PII is often the focus of privacy concerns, it is not the only source of re-identification of individuals. A study by Sweeney et al. (Sweeney, 2002) demonstrated that using just three attributes—postal code, date of birth, and sex—over 86% of U.S. citizens could be uniquely identified. These attributes, known as quasi-identifiers, may not reveal any information when used individually, but when combined, they can effectively re-identify individuals.
@@
Our work makes several important contributions:
@@
We provide our replication package  [^1] https://github.com/ConferencePurposes/log_privacy for future replication or extension. The rest of the paper is organized as follows. Section 2 presents a preliminary study and sets the context of our work. Section 3 analyzes data privacy regulations. Section 4 reviews existing literature and tools. Section 5 presents insights from our industry survey. Section 6 discusses the implications of our findings, and finally, Section 9 concludes our paper.
@@ 2.1. Approach
In our analysis, we examine 25 log datasets drawn from two well-known collections: LogHub (Zhu et al., 2023) and the Internet Traffic Archive (ITA) (Laboratory, 1998). These datasets provide a diverse range of logs, covering various domains and system types, which allows us to explore a broad spectrum of attributes and formats. Specifically, from the LogHub collection, we analyze all the datasets, including HDFS, Hadoop, Spark, Zookeeper, OpenStack, BGL, HPC, Thunderbird, Windows, Linux, Mac, Android, HealthApp, Apache, OpenSSH, and Proxifier. From the ITA collection, we consider Calgary, Saskatchewan, Boston, WorldCup98, NASA, ClarkNet, EPA, and SDSC. We also consider the YouTube dataset collected by Zink et al. (Michael Zink and Kurose, 2008). For detailed information about each dataset, please refer to the references.
To identify common attributes across different log datasets, we start by parsing each of the 25 log datasets using Drain (He et al., 2017), a well-known log parsing tool. The goal is to capture log templates that represent the structure of the log entries. After obtaining these templates, we meticulously analyze them to identify all the potential attributes that a log statement might contain. Additionally, we try to find any information within these attributes that could be potentially considered sensitive.
@@ 4.1.1. Searching Research Databases
We conduct our literature search using two well-known databases: the IEEE Xplore Digital Library [^2] https://ieeexplore.ieee.org/ and the ACM Digital Library [^3] https://dl.acm.org/. Our objective is to find papers from both conference proceedings and journals. To identify articles that focus on defining privacy in software logs or perform log anonymization, we employ the following search query:
@@ 4.1.3. Forward and Backward Snowballing
To ensure our study’s comprehensiveness, we conduct both forward and backward snowballing on the initially selected 22 papers. In this phase, we examine the papers that have cited these selected articles (i.e., forward snowballing) as well as the references cited within them (i.e., backward snowballing). For each of these papers, we apply our article selection criteria (detailed in Section 4.1.2) to determine if they meet our inclusion and exclusion criteria. This process results in the identification of 36 new articles. By combining these with the articles initially found through our database search, we compile a total of 58 papers that specifically address privacy in software logs.
@@ 4.1.4. Information Extraction
In the last step, we carefully examine how each paper defines privacy within the context of log data and identify the specific attributes they consider sensitive.
@@ 5. Industry Practices
In this section, we design a survey to understand industry professionals’ perspectives on software log privacy. The survey explores which attributes in software logs are considered sensitive by those involved in log management and anonymization. By gathering insights from experts, we aim to complement our findings from log analysis, regulatory review, and literature research, providing a comprehensive overview of prioritized attributes and handling approaches for privacy protection.
@@ 5.1. Survey Design
In designing the survey, we use a variety of question types, including yes/no questions, multiple-choice questions, Likert-scale questions, and open-ended questions, to capture both quantitative and qualitative insights from participants. Using these different types of questions, we aim to gather comprehensive data that includes both statistical and descriptive insights, allowing us to analyze not just what professionals think but also how they justify their opinions and approaches.
@@ 5.2. Participant Selection Strategy
According to Kitchenham et al. (Kitchenham and Pfleeger, 2008), it is crucial that the target population is capable of addressing the research questions effectively. To identify the relevant audience for software log privacy, we concentrate on professionals involved in log management, log analysis, and data privacy. We select survey participants using two methods. The first method involves searching for specific roles and keywords on LinkedIn [^4] https://www.linkedin.com/, a popular professional networking platform. We use four keywords: (1) privacy engineer, (2) data privacy officer, (3) IT manager, and (4) IT security engineer. These keywords help us filter professionals likely to have direct experience with software logs and data privacy. The second method involves distributing the survey link through the authors’ network to reach industry partners and professionals. Due to the broad distribution, we lack the exact count of recipients, preventing us from calculating a precise participation rate.
@@ 5.3. Survey Questions
Multiple-choise questions:
@@ 5.5. Survey Results
Table 5. The sensitive log attributes from industry perspective
@@ 6.1. What Should be Anonymized in Software Logs?
From an industry perspective, IP addresses, MAC addresses, hostnames, file paths, and various IDs are recognized as the most sensitive attributes in software logs. These attributes are frequently cited by experts and commonly appear in publicly available logs. For example, 86% of survey respondents view IP addresses as sensitive, and they are found in 80% of software logs. while IP addresses have been well-researched, other attributes are less studied. For instance, although 52% of industry participants consider file paths sensitive, and they are present in 72% of analyzed logs, they are rarely discussed in the literature, with only 7% of reviewed articles addressing them.
@@
6.2. Research Gaps and Future Directions.
@@ 7. Threats to Validity
Construct Validity. The survey questions may not capture all aspects of privacy concerns in software logs and could concentrate on certain log types or attributes. To mitigate this, we first conducted a preliminary study to identify the most common attributes in software logs and designed our survey accordingly. Additionally, the parsing tool used in our log analysis might have accuracy limitations, and some log attributes may have been missed. However, by combining insights from log analysis, data regulations, research articles, and industry experts, we aimed to achieve a comprehensive view of sensitive attributes in software logs.
@@ 8. Conclusion
We conduct a comprehensive analysis of privacy in software logs by examining the issue from multiple perspectives: regulatory frameworks, research literature, and industry practices. Our analysis of 25 publicly available log datasets identifies the most common log attributes that may require anonymization, such as IP addresses, MAC addresses, and hostnames. Our review of privacy regulations, including GDPR and CCPA, highlights the legal obligations to protect such sensitive data. Additionally, our systematic literature review reveals that while IP addresses and timestamps are commonly studied, other attributes, such as MAC addresses and file paths, are underexplored. Insights from a survey of 45 industry professionals further emphasize the challenges in balancing privacy and utility in log data, as well as in dealing with computational costs and processing time.
@@ 9. Data Availability
To support future replication or extension of our research, we have made our replication package available. This includes the list of log datasets analyzed, a list of reviewed articles, and detailed survey questions. Survey responses are also available upon request. The package can be accessed at https://github.com/ConferencePurposes/log_privacy.

### v2

# Protecting Privacy in Software Logs: What Should Be Anonymized?

@@ Abstract.
copyright: acmlicenseddoi: 10.1145/3715779journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE060publicationmonth: 7ccs: Software and its engineering Software safetyccs: Security and privacy Privacy protections
@@ 1. Introduction
Software logs are record statements used by software developers to capture valuable runtime information about software systems. They are generated by logging statements inserted into the source code, which produce execution logs during runtime (Barik et al., 2016; Li et al., 2020). These logs are then either appended to log files for later analysis or output to the standard console for immediate monitoring. Logging is crucial for tasks such as failure diagnosis and anomaly detection (Cinque et al., 2012; Batoun et al., 2024; Qin et al., 2024; Aghili et al., 2023). For instance, logs are often the only resource available for diagnosing field failures (Yuan et al., 2010). By recording different events and activities within a software system, logs help developers and other software practitioners manage and maintain the health and performance of their systems effectively.
As the number of log statements in software systems increases, the likelihood of capturing sensitive information within logs also rises, making privacy preservation more critical (Jain et al., 2016). Protecting privacy is essential not only to comply with regulatory requirements but also to maintain user trust and prevent potential harm that could arise from data leaks. A significant concern in software log management and sharing is the exposure of Personally Identifiable Information (PII). PII refers to any data that can be used to identify an individual, either directly, such as names and email addresses, or indirectly, such as device identifiers. While PII is often the focus of privacy concerns, it is not the only source of re-identification of individuals. A study by Sweeney et al. (Sweeney, 2002) demonstrated that using just three attributes—postal code, date of birth, and sex—over 86% of U.S. citizens could be uniquely identified. These attributes, known as quasi-identifiers, may not reveal any information when used individually, but when combined, they can effectively re-identify individuals.
@@
Our work makes several contributions:
@@
The rest of the paper is organized as follows. Section 2 analyzes the common attributes in software logs. Section 3 explores data privacy regulations. Section 4 reviews relevant literature and tools on log privacy. Section 5 presents insights from our industry survey. Section 6 discusses the implications of our findings, while section 7 addresses threats to validity. Finally, Section 8 concludes the paper, and Section 9 provides our replication package for future studies.
@@ 2.1. Approach
In our analysis, we examine 25 log datasets drawn from two well-known collections: LogHub (Zhu et al., 2023) and the Internet Traffic Archive (ITA) (Laboratory, 1998). These datasets provide a wide range of logs from different domains and system types, enabling us to investigate various attributes and formats. Specifically, we analyze all datasets from the LogHub collection, which includes logs from distributed systems (HDFS, Hadoop, Spark, Zookeeper, OpenStack), supercomputers (BGL, HPC, Thunderbird), operating systems (Windows, Linux, Mac), mobile systems (Android, HealthApp), server applications (Apache, OpenSSH), and standalone software (Proxifier). Additionally, from the ITA collection, we focus on web application logs, including datasets from Calgary, Saskatchewan, Boston, WorldCup98, NASA, ClarkNet, EPA, and SDSC. We also incorporate the YouTube dataset collected by Zink et al. (Michael Zink and Kurose, 2008). For detailed information about each dataset, please refer to the references.
To identify common attributes across different log datasets, we start by parsing each of the 25 log datasets using Drain (He et al., 2017), a well-known log parsing tool. Specifically, we analyze a sample of 2,000 log lines per dataset to extract log templates. We identify up to 350 unique log templates per dataset and then examine a subset of example log lines for each template to determine potential attributes. These attributes correspond to variables in the log structure, which Drain marks with an asterisk (*). To assess potential sensitive attributes, we take an inclusive approach, collecting any attribute that could pose a vulnerability. We later refine this by considering different perspectives on sensitivity (see Sections 3, 4, and 5) and ultimately defining our own criteria (see Section 6).
@@ 4.1.1. Searching Research Databases
We conduct our literature search using two well-known databases: the IEEE Xplore Digital Library [^1] https://ieeexplore.ieee.org/ and the ACM Digital Library. [^2] https://dl.acm.org/ Our objective is to find papers from both conference proceedings and journals. To identify articles that focus on defining privacy in software logs or perform log anonymization, we employ the following search query:
@@ 4.1.3. Forward and Backward Snowballing
To ensure thoroughness in our study, we conduct both forward and backward snowballing on the initially selected 22 papers. In this phase, we examine the papers that have cited these selected articles (i.e., forward snowballing) as well as the references cited within them (i.e., backward snowballing). For each of these papers, we apply our article selection criteria (detailed in Section 4.1.2) to determine if they meet our inclusion and exclusion criteria. This process results in the identification of 36 new articles. By combining these with the articles initially found through our database search, we compile a total of 58 papers that specifically address privacy in software logs.
@@ 4.1.4. Information Extraction
In the final step, we analyze how each paper defines privacy in the context of log data. We review their approach sections to determine which attributes are considered sensitive.
@@ 5. Industry Practices
In this section, we design a survey to understand industry professionals’ perspectives on software log privacy. The survey explores which attributes in software logs are considered sensitive by those involved in log management and anonymization. By gathering insights from experts, we aim to complement our findings from log analysis, regulatory review, and literature research, providing an overview of prioritized attributes and handling approaches for privacy protection.
@@ 5.1. Survey Design
In designing the survey, we use a variety of question types, including yes/no questions, multiple-choice questions, Likert-scale questions, and open-ended questions, to capture both quantitative and qualitative insights from participants. Using these different types of questions, we aim to gather diverse data that includes both statistical and descriptive insights, allowing us to analyze not just what professionals think but also how they justify their opinions and approaches.
@@ 5.2. Participant Selection Strategy
According to Kitchenham et al. (Kitchenham and Pfleeger, 2008), it is crucial that the target population is capable of addressing the research questions effectively. To identify the relevant audience for software log privacy, we concentrate on professionals involved in log management, log analysis, and data privacy. We select survey participants using two methods. The first method involves searching for specific roles and keywords on LinkedIn, [^3] https://www.linkedin.com/ a popular professional networking platform. We use four keywords: (1) privacy engineer, (2) data privacy officer, (3) IT manager, and (4) IT security engineer. These keywords help us filter professionals likely to have direct experience with software logs and data privacy. The second method involves distributing the survey link through the authors’ network to reach industry partners and professionals. Due to the broad distribution, we lack the exact count of recipients, preventing us from calculating a precise participation rate.
@@ 5.3. Survey Questions
Multiple-choice questions:
@@ 5.5. Survey Results
Table 5. Sensitive log attributes from industry perspective
@@ 6.1. What Should be Anonymized in Software Logs?
Figure 2 presents the sensitive log attributes from three different perspectives: privacy regulations, articles and tools, and industry practices. From an industry perspective, IP addresses, MAC addresses, hostnames, file paths, and various IDs are recognized as the most sensitive attributes in software logs. These attributes are frequently cited by experts and commonly appear in publicly available logs. For example, 86% of survey respondents view IP addresses as sensitive, and they are found in 80% of software logs. While IP addresses have been well-researched, other attributes are less studied. For instance, although 52% of industry participants consider file paths sensitive, and they are present in 72% of analyzed logs, they are rarely discussed in the literature, with only 7% of reviewed articles addressing them.
@@
Figure 2. Sensitive log attributes from different perspectives
6.2. Research Gaps and Future Directions
@@
  4. (4)
    Examining privacy challenges in small-sized companies vs. larger organizations. Future research could explore whether data privacy practitioners in small and medium-sized enterprises encounter unique challenges compared to those in larger organizations. Small companies often have fewer resources, less regulatory oversight, and limited expertise in log data privacy, which may impact their ability to implement effective anonymization and compliance measures. Investigating these differences could help tailor privacy solutions and regulatory frameworks to better support organizations of varying sizes.
  5. (5)
    Comparing stated privacy policies with actual log data collection. Future research could investigate the gap between what companies claim in their privacy policies and the log data they actually collect. Analyzing discrepancies between stated commitments and real-world practices could provide insights into transparency issues and help develop stronger regulatory guidelines for log privacy.
@@ 7. Threats to Validity
Construct Validity. The survey questions may not capture all aspects of privacy concerns in software logs and could concentrate on certain log types or attributes. To mitigate this, we first conducted a preliminary study to identify the most common attributes in software logs and designed our survey accordingly. Additionally, the parsing tool used in our log analysis might have accuracy limitations, and some log attributes may have been missed. However, by combining insights from log analysis, data regulations, research articles, and industry experts, we aimed to achieve a well-rounded view of sensitive attributes in software logs.
@@ 8. Conclusion
We conduct an in-depth analysis of privacy in software logs by examining the issue from multiple perspectives: regulatory frameworks, research literature, and industry practices. Our analysis of 25 publicly available log datasets identifies the most common log attributes that may require anonymization, such as IP addresses, MAC addresses, and hostnames. Our review of privacy regulations, including GDPR and CCPA, highlights the legal obligations to protect such sensitive data. Additionally, our systematic literature review reveals that while IP addresses and timestamps are commonly studied, other attributes, such as MAC addresses and file paths, are underexplored. Insights from a survey of 45 industry professionals further emphasize the challenges in balancing privacy and utility in log data, as well as in dealing with computational costs and processing time.
@@ 9. Data Availability
To support future replication or extension of our research, we have made our replication package available. This includes the list of log datasets analyzed, the list of reviewed articles, detailed survey questionnaire, and the survey responses. The package can be accessed at https://github.com/mooselab/log_privacy.
Acknowledgments
We would like to gratefully acknowledge the Natural Sciences and Engineering Research Council of Canada (NSERC, #RGPIN-2021-03900) and the Fonds de recherche du Québec – Nature et technologies (FRQNT, #326866) for their funding support for this work.